DriveTrain, Inc. instituted a new process in October 2020. During October, 10,600 units were started in Department A. Of the units started, 8,150 were transferred to Department B, and 2,450 remained in Work-in-Process at October 31, 2020. The Work-in-Process at October 31, 2020, was 100% complete as to material costs and 50% complete as to conversion costs. Material costs of $28,620 and conversion costs of $37,500 were charged to Department A in October. What were the total costs transferred to Department B assuming Department A uses weighted-

average process costing?
Answer #1

Material cost per unit = total material cost/(total units) = 28,620/(8.150+2,450) = 2.7

Conversion cost per unit = total conversion cost/(units transferred + 50% of WIP) = 37,500/(8,150+50%*2,450) = 4

So cost transferred to B = (material cost per unit + conversion cost per unit)*no of units = (2.7+4)*8,150 = 54,605

I hope it is useful to u if u have any doubt plz comment and plz give me up thumb